Could you please explain this one to me. Instructions are:
Determine th slope and the y-intercept of the garph of each equation.
x + 7y = 6
7y = -x + 6 I know that I need to divide by 7
but I come out with y = -1/7 + 6/7 is this correct?
Carolyn

Answer by tutorcecilia(2152)

You are almost correct, except that you forgot to put "x" next to the value of the slope:
x + 7y = 6
7y=-x+6
7y/7=-x/7+6/7
y=-1/7x+6/7
So, the slope is -1/7 and the y-intercept is 6/7